OpenStudy (anonymous):

Let f(x) = x2 - 16. Find f-1(x)

OpenStudy (aum):

To find the inverse of f(x): 1. Replace f(x) with y: y = x^2 - 16 2. Switch 'x' and 'y': x = y^2 - 16 3. Solve for y: x + 16 = y^2. y = sqrt(x+16). 4. Replace y with f-inverse or f^(-1)(x): f^(-1)(x) = sqrt(x+16)

OpenStudy (solomonzelman):

I was thinking, \(\large\color{slate}{ \displaystyle y=x^2-16}\) \(\large\color{slate}{ \displaystyle 0=x^2+0x-(16-y)}\) \(\large\color{slate}{ \displaystyle x=\frac{-(0)\pm\sqrt{0^2-4(1)(16-y)}}{2(1)} }\) \(\large\color{slate}{ \displaystyle x=\frac{\pm\sqrt{4(1)(y-16)}}{2} }\) \(\large\color{slate}{ \displaystyle x=\pm\sqrt{(y-16)} }\) and in inverse you replace, so \(\large\color{slate}{ \displaystyle f^{-1}(x)=\pm\sqrt{(x-16)} }\)

OpenStudy (solomonzelman):

So I think the \(\large\color{slate}{ \displaystyle \pm }\) should be there.

OpenStudy (aum):

\[ \large\color{slate}{ \displaystyle f^{-1}(x)=\pm\sqrt{(x+16)} } \]